09 June 2012 (UT) 1st half SciOps Instrument Check Out

  • Doug and Juan Carlos started, using LBBG & IRTC, to check out a bug they found last night when testing the AO with TCS build 2012F. The bug was fixed in the afternoon, and their start of night tests confirmed it was fixed.

  • 20:30 Geno reconfigured for MODS.
    • Ran a spectrophot std acq script which had been reorganized to take the field image, pause for modsAlign and then take a thru-slit confirmatory image. Worked fine.
#
# MODS Target Acquisition File
#
# Modifications: 2012 Mar 20 [rwp/osu]
#
# GD140 at PA=180 and the Spectrophotometric 5-arcsec slit
#
# There are no guide stars in the patrol field for PA=0
#
# Takes a single thru-slit image, if we are far off, we can
# follow with an acq snap
#

ACQMODE LongSlit

Archive:
  Partner CALIBRATION
  PropID CALIBRATION
  PI_Name All

Target:
  OBJNAME GD140 PA180
  OBJCOORDS 11:37:05.613 +29:47:59.33 -148.2 -5.3 2000
  GUINAME GStar
  GUICOORDS 11:37:07.524 +29:50:45.49
  POSANGLE 180
  AGWFILT Clear
  PRESET ACTIVE

Instrument:
  SLITMASK LS60x5
  ACQCamera Red
  ACQFilter r_sdss
  ACQexptime 2
  ACQROI 1Kx1K

Acquire:
  AcqGO
  print Run modsAlign -r to put the star at the slit reference position
  PAUSE
  SlitGO

end
  • Star was far off slit - proper motion was not initially included. It now is there, but still the star is not centered (wrong coordinates? wrong epoch?)
  • Tested new version of modsAlign, version 1.1.17 (current version is v1.15) v1.1.17 includes:
    • accounts for field distortion
    • outputs a file called mods_lastOffset which contains, verbatim, the offset command that was output by modsAlign and used to center the star on the slit. This could be incorporated in an acquisition script to be used when we need to reacquire for some reason, e.g. guiding lost.

  • mods1r images, starting with 9, are noisy (see 9,10,11,12). We cntl-C'd the script.
  • noticed that there was NO vacuum reading (Housekeeping - MODS1 environmental monitor). Blank.
  • Stopped IC and restarted... noticed vacuum N/A
  • Wanted then to power cycle the HEB (but accidentally powered off the IEB... this cut the power to all of red channel mechanisms).
    • recovery from IEB power off:
      • shut down all MODS services (following shutdown procedure)
      • restart all MODS services (following this time startup procedure)
      • modsColdStart
  • ran a sieveSnap: mods1r image number 13 is bad --- central row looks smeared horizontally. Redid the sieveSnap #14 and this is normal. (1536,1562).
  • ran grlamps.cal
  • From 14 on, mods1r images look ok.

-- OlgaKuhn - 09 Jun 2012
Topic revision: r2 - 09 Jun 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback